Determine the dimensions of q when q=1/2pv^2 when p =density and v=velocity
1
Expert's answer
2020-04-01T10:12:48-0400
"q=\frac{1}{2}\rho v^2.""[q]=[\rho]\times [v]^2""=\rm kg/m^3\times (m/s)^2=kg/(m\cdot s^2)=N/m^2=Pa."